Archive for May 2017

Apiary Record 28 May 2017

Weather cooler, following very warm spell. Rain and storms forecast. Objectives of the session were to make up an Observation Hive for display at Harting Festivities on 29 May and for new Beekeepers to check progress of their nucs. Hives 3 and 8 inspected, with varroa checks on all hives.

Hive #1 – not checked. Varroa = >100.
Hive #2 – Wednesday inspection.
Hive #3 – inspected by Richard and Liz. Double brood with 3 supers. Queen not seen. Knocked down all but two Queen cells – one capped and one open. Varroa count = 17.
Hive #4 – not checked. Varroa count = 14.
Hive #5 – not checked. Next inspection due 11 June (leaving for 3 weeks). Varroa count = 13.
Hive #6 – Wednesday inspection.
Hive #7 – not checked. Varroa count = 5.
Hive #8 – inspected by David P. No Queen cells drawn since last week. Brood emerging. Eggs seen in top brood box but queen may not be mated. Swapped brood boxes around so queen can fly for mating. Needs frame of eggs. High varroa count >130.

Apiary Record 24 May 2017

Hives 2, 2A and 6 inspected by Liz. Weather warm and sunny.

Hive #2 – Varroa count = 4 (board inserted 3 days ago). Good temper.
Hive #2A – No varroa board. Dounble brood. 1 emerged queen cell with flap still present. Mesh floor had been placed upside down – now corrected. Good temper.
Hive #6 – Varroa count 60+ (board inserted 3 days ago). Lots of larvae but no eggs seen. Good temper.

Apiary Records 21 May 2017

Weather warm and sunny, 20 degrees. All hives inspected except 2 & 6 which are reserved for mid-week inspections. Varroa boards added.
Hive #1 – inspected by Richard & Liz. 1 sealed queen cell. No eggs or larvae. Frame with QC moved to #5. Top super has frames to extract.
Hive #3 – inspected by Richard & Liz. Queen found and marked white. Removed to Nuc. 4 frames taken to extract. Evidence of Deformed Wing Virus.
Hive #4 – inspected by Anne. Should have been left for 4 weeks with next inspection on 28 May. Leave for a further 3 weeks.
Hive #5 – inspected by Mel. No evidence of Queen. Many supercedure cells (frame 9 had 9 cells). Removed additional half brood and placed above Queen Excluder. Leave for 3 weeks.
Hive # 7 – inspected by Graham. Brood in all stages. 2 deep frames of brood removed for Nuc and replaced with indrawn foundation. If Queen not found next week, consider putting bees through QX. Possible use 2 frames for show.
Hive #8 – top brood box had frame of eggs added from # 6 as brood too old. Bottom box not drawn so possibly no Queen.

Association Library up and running!

The library book list, edited and updated, is now on the ‘Library’ page on the ‘About us’ Page on this website.  The books will be available to members at all Association Meetings, both Apiary and Evening, and via this website with a link to the Librarian.  With old favourites and new purchases there is something for everyone.

Apiary Record Sunday 14th May

Breezy, 18C  David Parkinson making Nuc Hives and initiating queen rearing.

Hive #1  Inspected by Richard Williams.  Good temper.   14×12 +shallow.  2 supers.   4 of 5 QC’s removed.  Sealed brood only. Sufficient space and stores.

Hive #3  Inspected by Liz Hawes.  Good temper.  2 deep.  3 supers.  Queen seen, not marked. 2 QC’s removed.  Sufficient space .

Hive #5  Inspected by Anne-Chantal Ballard.  Good temper. 14×12.  2 supers.  9 brood, eggs, 4 QC’s. Shallow added to brood.  QC’s & 1 framebrood removed to nucs.


Apiary Record 10th May

Sunny 20C

Hive #2  Inspected by Tom Blackburn  Artificial Swarm performed to create Hive #2a.

Hive #2 1 deep 1 super , queen marked blue, 2brood, 2 stores.

Hive #2a  2 deep, 2 sealed QC’s.Sufficient stores in deep

Hive #6  Inspected by Tom Blackburn.  2 deep, 6 & 7.5 frames brood 3 brood frames stores;  2 supers with limited stores.

Apiary Record Sunday 7th May

Sunny 17C .

Hive #1 Inspected by Liz Hawes.  14×12 + shallow.  2 supers.  Sufficient room & stores . Good Temper.  Queen not seen, swarm suspected. QC’s: 1 hatched, 2 torn down, 3 small, 2 good remaining.

Hive # 3  Inspected by Rowan Roberts.  2 deeps. 2 supers.  Sufficient room & stores . moderate Temper.  Queen not seen. 1 Queen cell removed to Nuc. hive, others destroyed.

Hive #4  Not inspected

Hive #5   Inspected by Rowan Roberts.  14×12. 2 supers.  Sufficient room & stores . moderate Temper.  Queen seen. 9 frames brood.  Bias.  Drone trap moved to brood edge.

Hive #7  Not inspected

Hive #8  Inspected by Phil Gurney. 1deep + shallow. 1 super.  good temper. 9 deep, 1 shallow brood, 1 torn down QC.  no eggs seen, young larvae.  Stores in brood only.

Hive # 9    Inspected by Phil Gurney. 1 deep . 1 super.  good temper.  No eggs, little larvae.  Poor condition, Sac brood.  Little stores, little feed taken.  Graham Rowden took Decision to destroy



Apiary Record 3rd May

Cloudy, cold & turning wet.  Quick inspection of Hive#2; very brief of hive#6 Tom Blackburn leading.

Hive #2    2 deeps with 4 brood in each.  Unsealed QC removed. 1 super with 9.5 stores.  Space.

Hive #6  Deteriorating weather, bad tempered bees.  Stores only inspected, 4 frames.


Apiary Meeting 30 April

Too cold for general inspections – 10 degrees. Demonstration of Queen marking using workers found in supers. Checks on nucs created the previous week.

Hive #4 – inspected by David P. 3 supercedure cells, likely to be emergency supercedures. Two cells cut from hive and pasted into Nuc 4A. Bees from some frames shaken into Nuc. Leave for 4 weeks. Next inspection 28 May.

Hive #3A – inspected by David P. Nuc created 23 April. Queenless.

Hive 4A – inspected by David P. Two supercedure cells pasted into Nuc from #4. Bees shaken into Nuc from some #4 frames.

Apiary Meeting 26 April

Brief inspections on Hives 2 & 6 only. Weather cold and wet.

Hive #2 – inspected by Liz & Rowan. Deep x 2. Brood on 4 frames. Stores on 5 frames. Some play cups. Stores in tip deep but not in super.

Hive #6 – inspected by Liz & Rowan. Deep x 2. Play cups seen in top deep. Only 5 frames inspected in bottom deep – hail storm prevented further checks.